Two cars leave an intersection. One car travels north; the other east. When the car travelling north had gone 12 miles, the distance between the cars was 8 miles more than the distance traveled by the car heading east. How far had the eastbound car traveled?
Answer by mananth(16946)

As you can see they drive at right angles
..
let east bound travel x miles.
north bound has traveled 12 miles by then.
The distance between them = x+8 ( hypotenuse)
..
pythagoras theorem
(x+8)^2=x^+12^2
x^2+16x+64=x^2+144
16x+64=144
add -64
16x+64-64=144-64
16x=80
x= 5 miles. distance traveled by East bound